suppose the population P(t) of Ongwediva town at any time t>0 satisfies the logic law
dP/dt=P/100-P^2/〖10〗^8
Where t is measured in years. Given that the population of this city was 100000 in 1980 determine the population as a function t>1980
1
Expert's answer
2020-12-02T13:15:38-0500

"\frac{dP}{dt}=\frac{P}{100}-\frac{P^2}{10^8}\\
10^8\frac{dP}{10^6P-P^2}=dt"

"10^8\int{\frac{dP}{10^6P-P^2}}=\int{dt}"

"100ln\frac{P}{P-10^6}=t+C_1"

"\frac{P}{P-10^6}=Ce^{\frac{t}{100}}"


Given that in the year 1980, t=0 and "x(0)=10^5" we get "C=-\frac{1}{9}"


Then, "\frac{P}{P-10^6}=-\frac{1}{9}e^{\frac{t}{100}}"

"P(t)=\frac{10^6e^{\frac{t}{100}}}{e^{\frac{t}{100}}+9}"